A yacht leaves port and sails 50km due east. It then changes direction and sails another 50km on bearing of 200 degrees. How far from the port is the yacht now?
1
Expert's answer
2012-08-28T10:26:46-0400
According the low of cosines, c^2 = a^2 + b^2 - 2*a*b*cos (gamma);
let a = b = 50 miles and gamma = 200 degrees.


c^2 = 50^2 + 50^2 - 2*50*50*cos 200 = 2500 + 2500 - 5000* (- sin 70) = 5000*(1+sin 70) ~ 5000*1.9397 ~ 9698.4631
c = 98.4808